n your new role as a data analyst for a large hospital chain, the CEO has asked you to explore patient outcomes at a newly built hospital (“Mountain View”) versus an existing hospital (“Greenside”). She is specifically interested in seeing whether patients who are seen for heart attacks are being readmitted to the hospital more often in the new facility. You looked at data from a random sample of visits to the two hospitals during the last year. Use the following table to answer questions 12-16.

Was Patient Readmitted to the Hospital within 30-days following treatment for heart attack?

Hospital Site Visited

Mountain View

Greenside

Total

Yes

54

16

70

No

705

171

876

Total

759

187

946


2. If you choose a person from the sample at random, what is the probability that the person was readmitted to the hospital within 30-days following treatment for heart attack?

3. If you choose a person from the survey at random, what is the probability that the person visited the new Mountain View hospital site?

4.If you choose a person from the survey at random, what is the probability that the person had visited the Greenside hospital and was not readmitted to the hospital within 30-days following treatment for heart attack?
Answer #1

Answer:

2) The probability that the person was readmitted to the hospital within 30-days following treatment for heart attack:

= 70/946 = 0.074 = 7.4%

3) The  probability that the person visited the new Mountain View hospital site:

= 759/946 = 0.8023 = 80.23%

4) The  probability that the person had visited the Green side hospital and was not readmitted to the hospital within 30-days following treatment for heart attack:

= 171/946 = 0.1808 = 18.08%